# Ballistic conductance with and without disorder in a boundary-driven XXZ spin chain

### v2

@@ 2 Model
We consider a classical chain of unit-length spins ${\boldsymbol{S}_{i}\in S^{2}}$ with boundary-driven Floquet dynamics. A single time-step $t\mapsto t+\delta t$ consists of three parts. Between times $t$ and $t+\delta t/2$, odd bonds $(1,2)$, $(3,4)$, …, $(L-1,L)$ evolve, followed by the even bonds between times $t+\delta t/2$ and $t+\delta t$. Finally, the boundary spins are reset, $\boldsymbol{S}_{1}=-\boldsymbol{\hat{z}}$, $\boldsymbol{S}_{L}=+\boldsymbol{\hat{z}}$ (cf. iSWAP($\pi/2$) gates of the quantum experiment [87]).
The bond dynamics is generated by the restriction of the XXZ Hamiltonian to the pair of sites being evolved,
@@
The anisotropy $\Delta$ is the principal tuning parameter. Below, we first consider a clean system, $J_{i}=1$, followed by the bond disordered case.
@@ 4.2 Stability of the ballistic regime
Again, we begin by considering the steady-states in the infinite-frequency limit, given by the consistency equations (7), but now with random couplings. We show that, in the easy-plane, so long as there is some lower bound to the coupling strengths (i.e., $W<1\Rightarrow J_{i}>0$), the ballistic regime survives in the sense that $G(L)\sim j_{\infty}\sim O(L^{0})$. This stability derives from the fact that the disordered consistency equations are fundamentally limited in the extent to which they can modify the NESS of the clean model (We discuss typical disordered steady states in App. C).
To establish that the ballistic regime survives, we need to show that there remains a finite, length-independent difference between $z_{L-1}$ and $z_{L}=1$, since, cf. Eq. (8), this yields a finite conductance $G(L)\sim 1-z_{L-1}$.
We conjecture that the minimum conductance $G(L)$ (as a function of the disorder realisation) is obtained from the couplings $J_{2}=...=J_{L-2}=1-W$, and that this, indeed, corresponds to a finite, length-independent lower bound on $G(L)$. We show this minimum conductance as a function of the disorder strength $W$, for various anisotropies $\Delta<1$, in Fig. 4(a).
4.3 Finite-frequency simulations with disorder
@@ 4.4 Subdiffusion at strong disorder
Finally, let us briefly address what happens when the disorder strength reaches $W=1$. The arguments of §4.2 no longer apply, and, as shown in Fig. 5(a), we find a subdiffusive scaling of the disorder average of the injected current, $\overline{j_{1}(t)}\sim t^{-2/3}$. We discuss a possible explanation for this observation in App. D.
@@ Appendix B Solution of the consistency equations
Appendix C Typical disordered steady-states
@@
Appendix D Subdiffusion at strong disorder
We turn our attention now to the transition away from the ballistic regime at strong disorder: the arguments of §4.2 do not apply if there is no nonzero lower bound to the coupling strengths, $J_{i}\geq J_{\mathrm{min}}>0$, and, as shown in Fig. 5(a), the exponent jumps from ballistic ($\alpha=1$) to subdiffusive ($\alpha=1/3$) behaviour at $W=1$.
@@ Appendix D Subdiffusion at strong disorder
Figure 9: Time-dependence of the disorder-averaged injected current $\overline{j_{1}(t)}$, for the power-law distributions of disorder, Eq. (D.1). The power-law asymptotes use the subdiffusive exponent $\alpha(\eta)$ predicted by Eq. (D.3). Again, $10^{4}$ realisations of disorder have been averaged over. $L=1024$.
@@
$\overline{\min\limits_{i=1,...,L}J_{i}}\sim L^{-\frac{1}{1-\eta}}.$ (D.2)
@@
$G(L)\sim L^{-\frac{1}{1-\eta}-1}\;\;\;\Rightarrow\;\;\;\alpha=\frac{1}{2+\frac{1}{1-\eta}},$ (D.3)

### v3

@@ 2 Model
We consider a classical chain of unit-length spins ${\boldsymbol{S}_{i}\in S^{2}}$ with boundary-driven Floquet dynamics. A single time-step $t\mapsto t+\delta t$ consists of three parts (cf. Fig. 1):
  • odd bonds $(1,2)$, $(3,4)$, …, $(L-1,L)$ evolve between times $t$ and $t+\delta t/2$
  • even bonds $(2,3)$, $(4,5)$, …, $(L-2,L-1)$ evolve between times $t+\delta t/2$ and $t+\delta t$
  • Finally, the boundary spins are reset, $\boldsymbol{S}_{1}=-\boldsymbol{\hat{z}}$, $\boldsymbol{S}_{L}=+\boldsymbol{\hat{z}}$ (cf. iSWAP($\pi/2$) gates of the quantum experiment [87]).
Within each half-timestep, the bond dynamics are generated by the restriction of the XXZ Hamiltonian to the pair of sites being evolved,
@@
That is, the equations of motion are
$\dot{S}_{i}^{\mu}=\epsilon^{\mu\nu\lambda}J_{i}^{\nu}S_{i+1}^{\nu}S_{i}^{\lambda},\;\;\;\;\;\dot{S}_{i+1}^{\mu}=\epsilon^{\mu\nu\lambda}J_{i}^{\nu}S_{i}^{\nu}S_{i+1}^{\lambda}\;\;\;\;\;(J_{i}^{x}=J_{i}^{y}=J_{i},\;\;\;J_{i}^{z}=J_{i}\Delta),$ (2)
for the bond $(i,i+1)$ being evolved. The anisotropy $\Delta$ is the principal tuning parameter. Below, we first consider a clean system, $J_{i}=1$, in §3, followed by the bond disordered case in §4.
@@ 3.4 Classical-quantum correspondence
Indeed, it is this similarity in the steady-states that appears to be responsible for the detailed correspondence between the quantum and classical cases. As further evidence for this, we show in App. C that moving closer to an integrable point has no effect on the asymptotic transport, so long as the NESS is the same.
@@ 4.2 Stability of the ballistic regime
Again, we begin by considering the steady-states in the infinite-frequency limit, given by the consistency equations (8), but now with random couplings. We show that, in the easy-plane, so long as there is some lower bound to the coupling strengths (i.e., $W<1\Rightarrow J_{i}>0$), the ballistic regime survives in the sense that $G(L)\sim j_{\infty}\sim O(L^{0})$. This stability derives from the fact that the disordered consistency equations are fundamentally limited in the extent to which they can modify the NESS of the clean model (We show examples of typical disordered steady states in App. D). In particular, the presence of disorder does not change the fact that the infinite-frequency steady states are monotonic, that $z_{i}\approx 0$ in the bulk; and there remains a finite difference between $z_{L-1}$ and $z_{L}=1$, which determines the transport properties.
To establish that the ballistic regime survives, we need to show that there remains a finite, length-independent difference between $z_{L-1}$ and $z_{L}=1$, for any configuration of disorder, since, cf. Eq. (9), this yields a finite conductance $G(L)\sim 1-z_{L-1}$.
We conjecture that the minimum conductance $G(L)$ (as a function of the disorder realisation) is obtained from the couplings $J_{2}=...=J_{L-2}=1-W$, and that this, indeed, corresponds to a finite, length-independent lower bound on $G(L)$. We show this minimum difference as a function of the disorder strength $W$, for various anisotropies $\Delta<1$, in Fig. 4(a).
4.3 Finite-frequency simulations with disorder
@@ 4.4 Subdiffusion at strong disorder
Finally, let us briefly address what happens when the disorder strength reaches $W=1$. The arguments of §4.2 no longer apply, and, as shown in Fig. 5(a), we find a subdiffusive scaling of the disorder average of the injected current, $\overline{j_{1}(t)}\sim t^{-2/3}$. We discuss a possible explanation for this observation in App. E.
@@ Appendix B Solution of the consistency equations
Appendix C Effect of integrability
We turn our attention now to the role of integrability in determining the transport regimes. We show in this appendix that the Ishimori chain, which is fully integrable, exhibits the same phenomenology in the long-time asymptotics.
We thus consider the corresponding Hamiltonian to generate the bond dynamics (cf. Eq. 1), though each timestep still consists of the same three parts: namely, that the odd bonds evolve over the first half-timestep; the even bonds evolve over the second half-timestep; and then the boundary spins are reset to $\boldsymbol{S}_{1}=-\boldsymbol{\hat{z}}$, $\boldsymbol{S}_{L}=+\boldsymbol{\hat{z}}$. The same initial state (5) as the main text is used.
The classical integrable Ishimori chain [6] has the Hamiltonian
$\mathcal{H}=-J\sum_{i}\log\left(1+\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+1}\right).$ (C.1)
To investigate the effect of proximity to an integrable model, we interpolate between the Ishimori chain and the classical Heisenberg chain with the following Hamiltonian:
$\mathcal{H}=-2J\gamma^{-1}\sum_{i}\log\left(1+\gamma\frac{\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+1}-1}{2}\right),$ (C.2)
which is the integrable Ishimori chain when $\gamma=1$, and the non-integrable classical Heisenberg chain when $\gamma\to 0$. We are only considering the clean, isotropic case in this appendix.
We show the time-dependence of the injected current (cf. Fig. 3(a)) for various values of $\gamma$ in Fig. 8. The short time dynamics does depend on $\gamma$, but in all cases, including the integrable case $\gamma=1$, the same decay $j_{1}(t)\sim t^{-2/3}$ is observed as for the classical Heisenberg chain shown in the main text, implying subdiffusive transport with $\alpha=1/3$. This is consistent with the picture developed in the main text, where the asymptotic transport can be deduced almost entirely from the NESS.
Indeed, the continuous-time equations of motion for the interpolating Hamiltonian are
$\dot{\boldsymbol{S}}_{i}=2\boldsymbol{S}_{i}\times\left(\frac{\boldsymbol{S}_{i-1}}{2-\gamma+\gamma\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i-1}}+\frac{\boldsymbol{S}_{i+1}}{2-\gamma+\gamma\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+1}}\right),$ (C.3)
and so the infinite-frequency consistency equations ($\dot{\boldsymbol{S}}_{i}=0)$ are
$\frac{2z_{i-1}\sqrt{1-z_{i}^{2}}-2z_{i}\sqrt{1-z_{i-1}^{2}}}{2-\gamma+\gamma\sqrt{1-z_{i}^{2}}\sqrt{1-z_{i-1}^{2}}+\gamma z_{i}z_{i-1}}+\frac{2z_{i+1}\sqrt{1-z_{i}^{2}}-2z_{i}\sqrt{1-z_{i+1}^{2}}}{2-\gamma+\gamma\sqrt{1-z_{i}^{2}}\sqrt{1-z_{i+1}^{2}}+\gamma z_{i}z_{i+1}}=0.$ (C.4)
For any $\gamma$, these are solved by the same infinite-frequency NESS,
$z_{n}=\sin\left(\frac{\pi}{L}\left(n-\frac{L+1}{2}\right)\right),$ (C.5)
which may be verified by direct substitution.
We note in passing that the short time dynamics do seem to more closely resemble the quantum case (where there is no delay time) near the integrable point. However, this is likely just an artefact of how the initial state was chosen – with a fixed rotation $\chi=0.01$ for the second spin (cf. Eq. (5)). In all these classical cases, starting with $\boldsymbol{S}_{2}=-\boldsymbol{\hat{z}}$ is an unstable steady state, so the rotation remains necessary (unlike the quantum case). But the energy of this state increases as we approach the integrable point (and, indeed, diverges for $\chi\to 0$, $\gamma\to 1$), which is presumably responsible for the reduced delay time.
Figure 8: Injected spin current $j_{1}(t)$ of the (clean) isotropic chain (C.2) for various values of the integrability parameter $\gamma$. In all cases, the same subdiffusive transport is observed, with $\alpha=1/3$; which is also the transport exponent in the isotropic case of the quantum experiment [87]. $L=1024$, $\delta t=\pi/8$.
Appendix D Typical disordered steady-states
@@
Appendix E Subdiffusion at strong disorder
Finally, let us consider the transition away from the ballistic regime at strong disorder: the arguments of §4.2 do not apply if there is no nonzero lower bound to the coupling strengths, $J_{i}\geq J_{\mathrm{min}}>0$, and, as shown in Fig. 5(a), the exponent jumps from ballistic ($\alpha=1$) to subdiffusive ($\alpha=1/3$) behaviour at $W=1$.
@@ Appendix E Subdiffusion at strong disorder
Figure 10: Time-dependence of the disorder-averaged injected current $\overline{j_{1}(t)}$, for the power-law distributions of disorder, Eq. (E.1). The power-law asymptotes use the subdiffusive exponent $\alpha(\eta)$ predicted by Eq. (E.3). Again, $10^{4}$ realisations of disorder have been averaged over. $L=1024$.
@@
$\overline{\min\limits_{i=1,...,L}J_{i}}\sim L^{-\frac{1}{1-\eta}}.$ (E.2)
@@
$G(L)\sim L^{-\frac{1}{1-\eta}-1}\;\;\;\Rightarrow\;\;\;\alpha=\frac{1}{2+\frac{1}{1-\eta}},$ (E.3)